OpenStudy (anonymous):

solve the inequality 6(3x-1)<=8(2x-8)

OpenStudy (anonymous):

\[6(3x-1)\le8(2x-8)\] As you multiply, divide, and add. Remember that whatever you do, as long as you do the same thing to both sides you are ok so to start distribute the 6 in 6(3x-1) and get \[18x-6\le8(2x-8)\]

OpenStudy (anonymous):

then distribute the 8 in 8(2x-8) and get \[18x-6 \le 16x-64\]

OpenStudy (anonymous):

subtract 16x from both sides and get \[2x-6\le-64\]

OpenStudy (anonymous):

then add 6 to both sides and get \[2x \le -58\]

OpenStudy (anonymous):

then divide by 2 and get \[x \le -29\] that's your final answer.

OpenStudy (anonymous):

Thank you so much